OpenStudy (anonymous):

Find the difference between 10 and 2 9th's - 7 and two thirds?

OpenStudy (anonymous):

\[10 \frac{ 2 }{ 9 }- 7 \frac{ 2 }{ 3 } =\] Do you know how to proceed from there?

OpenStudy (anonymous):

No, not exactly...I'm sorry.

OpenStudy (anonymous):

The first step is to convert each of the two fractions into improper fractions. To do this you need to multiply the denominator by the whole number, and add it to the numerator. \[10\frac{ 2 }{ 9 } \rightarrow \frac{ 92 }{ 9 }\] So you try to do the other fraction.

OpenStudy (anonymous):

Are you stuck?

OpenStudy (anonymous):

Very... But I finished that question, I'm caught on a new one. I dunno how to answer it. Please respond quickly, the test is timed and I have another to do after this one.